A 80kg uniform beam is 10m long and supported horizontally at one end A and at point C which us 3m from the other end B.

What is the maximum load ( in Newtons) that can be placed from B before the beam tips? Hint: the beam will start to tip when the reaction force at point point.

Once you have the maximum load calculate the reaction force at point C.
1
Expert's answer
2017-10-25T15:16:07-0400
Forces in equilibrium:
mg+W=R_A+R_C
R_A=0
Moments in equilibrium:
mg(10/2-3)=W(3)
mg(2)=W(3)
The maximum load:
W=2/3 mg=2/3 80(10)=533 N.
The reaction force at point C:
R_C=mg+W=mg+2/3 mg=5/3 mg=5/3 80(10)=1333 N.